riable cost per unit is od are $750,000. Ellis variable costs increase by 20% and fixed cost The selling price of one unit oCould you solve clearly? Thanks a lot!!
Answer #1

Break even point = Fixed cost/Contribution margin per unit

Current break even point = 150,000/(4.50-1.50)

= 150,000/3 = 50,000 units

.

New variable cost = 1.50+20% increase = 1.80

New fixed cost = 150,000-5% decrease = 142,500

New break even point = 142,500/(4.50-1.80)

= 142,500/2.70 = 52,778 units

.

So, these changes increased Breakeven point by 2,778 units (52,778 - 50,000)

Option C is the answer